OpenStudy (medina13):

PLEASE HELP !!!! A. How many rational roots does x^4+3x^2+2=0 have? Question options: 2 4 1 0

OpenStudy (medina13):

Yeah obvi. i dont no how

OpenStudy (medina13):

How do i find the factors ?

OpenStudy (medina13):

numbers that can multiply together to get another number

OpenStudy (medina13):

2x1=2 ?

OpenStudy (medina13):

So the answer is 1 or 2 ?

OpenStudy (medina13):

ummmm no ?

OpenStudy (medina13):

no

OpenStudy (zzr0ck3r):

use rational root theorem rational roots will be of the form \(\frac{\pm1}{1,2}\text{ so }\{-1/2,-1,1,1/2\}\) are the only possible rational roots

OpenStudy (medina13):

I dont know that

OpenStudy (medina13):

Of course lol

OpenStudy (zzr0ck3r):

sure it does...well when someone asks for "rational roots" of a degree 4 polynomial I assume they are not meant to factor it... there is a theorem for it so I let the user know about it...you know...because its called the rational root theorem and he asked about rational roots.

OpenStudy (medina13):

3

OpenStudy (zzr0ck3r):

So in short, no I am not kidding you, and in fact I was not commenting to your post.

OpenStudy (zzr0ck3r):

correct and after checking the 4 possible ones, there are none and we are done.

OpenStudy (zzr0ck3r):

which is SUPER easy, hence the theorem...you know....math

OpenStudy (zzr0ck3r):

grow up

OpenStudy (medina13):

Math is the worst subject for me so i need all the help i can get

OpenStudy (medina13):

Anyone there lol @zzr0ck3r @Mertsj

OpenStudy (medina13):

Now im confused on what to do .-.

OpenStudy (zarkon):

\[ x^4+3x^2+2\ge 2\]

OpenStudy (medina13):

And what is that !

OpenStudy (anthonyym):

You can do this quickly with a graphing calculator. Try desmos. Type in the function, then how ever many times the graph crosses the x-axis is how many rational roots there are.

OpenStudy (zzr0ck3r):

I am sorry he\she left you. I was simply trying to show you another way and was not meaning to interrupt. The point was that you have the information. I am not sure where you guys were heading. But I will explain the solution I was referring to. You have a polynomial and there is a theorem that says that if \(\pm\dfrac{p}{q}\) is a rational root then p is a factor of the leading coefficient and q is a factor of the constant term. 1=1*1=-1(-1) 2=1*2=-1*-2 So the possible roots are \(-1,1,-1/2,1/2\) and you can easily check that none of these are roots. So it has none.

OpenStudy (medina13):

oh okay ! well that is alot easier way. thank you so muchh :)

OpenStudy (zarkon):

you might want to look at that again @zzr0ck3r

OpenStudy (zarkon):

Also, my way is the quickest ;) \(x^4\ge 0\), \(3x^2\ge 0\) thus \[x^4+3x^2+2\ge2>0\] thus it has no real zeros...hence no rational zeros

OpenStudy (zarkon):

for the rational root theorem p is a factor of the constant term and q is a factor of the leading term.

OpenStudy (medina13):

Okay so i just need to no that right and easy way lol

OpenStudy (zzr0ck3r):

woops sorry I had it upside down. ....q is a factor of the leading coefficient and p is a factor of the constant term..... {1,-1,2,-2}

OpenStudy (zzr0ck3r):

I now see you cleared that up and I am being redundant :)

OpenStudy (zzr0ck3r):

zarkons way FTW so far imo. Takes only basic middle school math.

OpenStudy (medina13):

wait so whats going on ?